OpenStudy (anonymous):

Find the first five terms of the given recursively defined sequence. A[n]= 1/ 1+ A[n]-1 and A1=2 A[2]= A[3]= A[4]= A[5]=

OpenStudy (anonymous):

is this \[a_n=\frac{1}{1+a_{n-1}}\]?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok so first we fine \[a_2\] via \[a_2=\frac{1}{1+a_1}=\frac{1}{1+2}=\frac{1}{3}\]

OpenStudy (anonymous):

Why don't we subtract 2-1= 1/2

OpenStudy (anonymous):

then repeat the process to find \[a_3=\frac{1}{1+a_2}=\frac{1}{1+\frac{1}{3}}=\frac{1}{\frac{4}{3}}=\frac{3}{4}\]

OpenStudy (anonymous):

the subscript does not mean subtract the term. the subscript is like saying "first term" "second term" "third term" and so on

OpenStudy (anonymous):

so this statement \[a_n=\frac{1}{1+a_{n-1}}\] means the "nth" term is found by taking one over one plus the previous term

OpenStudy (anonymous):

Oh ok...

OpenStudy (anonymous):

so a4=7/4 and a5= 11/4

OpenStudy (anonymous):

\[a_4=\frac{1}{1+a_3}=\frac{1}{1+\frac{3}{4}}=\frac{1}{\frac{7}{4}}=\frac{4}{7}\] right

OpenStudy (anonymous):

but that is not what i get for the next term

OpenStudy (anonymous):

i get \[a_5=\frac{1}{1+a_4}=\frac{1}{1+\frac{4}{7}}=\frac{1}{\frac{11}{7}}=\frac{7}{11}\]

OpenStudy (anonymous):

why is not what i had?

OpenStudy (anonymous):

you had 11/4 but my arithmetic gave 7/11